OpenStudy (saylilbaby):

Suppose a mother carries two recessive genes for freckles (rr) and a father carries one recessive gene for freckles and one dominant gene for clear facial skin (Rr). What is the likelihood that their first child will have freckles? A) 1 2 B) 1 4 C) 3 4 D) 4 4 B??? @Ahsome

OpenStudy (ahsome):

The best way is to draw a table showing both genes, and see the posibilities. There are two Rr, and two rr. Since we want to know what is the probability of clear skin, and that clear skin is from Rr or RR. Then the probability is: \[\text{Probability}=\frac{\text{Number Wanted}}{\text{Total Posibilities}}\]\[\text{Probability}=\frac{2}{4}\]\[\text{Probability}=\frac{1}{2}\]